Date:
Wednesday, September 30, 2015
Telescope:
CTIO Blanco 4-m
Submitter:
Suk Sien Tie
Submitter Email:
Observer Support:
Hernan Tirado
Night Assistant:
Alberto Alvarez
Problems:
1. Please provide decaf coffee in the control room kitchen.
2. The women bathroom are running out of toilet papers and sanitary pads.
3. Please fix or replace LG media hub and speaker on observer 1.
Program:
Proposal:
2012B-0001
Target of Oportunity:
No
Time Observed:
3.0
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
7.0
Time lost other:
0.0
Description:
Dark Energy Survey
Primary Investigator:
Josh Frieman
Email:
Institution:
Fermilab
Instrument:
CTIO 4m DECam
Organization:
NOAO
Total Program Time:
10
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2015-09-30 | 3.0 | 0.0 | 0.0 | 7.0 | 0.0 | 10 |
*Observing Team*
_DES Observers:_ Cristina Furlanetto, Suk Sien Tie, David James
_Telops:_ Hernan Tirado, Alberto Alvarez
_Remote Support:_ ---
*Observing Plan*
# standard bias and dome flat sequence
# evening standard stars
# run ObsTac queue
# end of night calibrations
*Conditions*
# weather in afternoon: 4pm (local): Heavy cirrus coming in from south west.
# weather after sunset: 8pm (local): Heavy clouds still hanging on the south west; Rasicam showing the clouds creeping in. Other parts of the sky are partially cloudy. Non-photometric. Standards taken anyway.
# weather at 11:30pm (local): Heavy clouds (can barely see any stars in DECam images) and concern over precipitation (forecast to rain tomorrow morning). Closing telescope at 11:33pm.
# weather at 2am (local): Overcast - can't even see the moon!
*Accomplishments*
# Afternoon: zeros done, (ugrizYVR) dome flats done, checked, OK
# Evening: Beginning of night standards are SDSSJ2300-000 (airmass 1.8), SDSSJ2140-000 (airmass 1.3), and E8-A (airmass 1.0).
# aTmCam and GPS Monitor Checked.
# ObsTac ran until 11:30pm local time.
# Survey Exposures 480580-480650 (g, r, i, and some z bands).
# Did DECal after closing telescope.
*Problems tagged as Synopsis_Problem*
# None
*Minor Problems and Relevant Information*
# Minor issue with SISPI one hour into flat-fielding where number of image-builders down to 0 and dts > 300s. Fixed this by stopping exposures to free up image-builders.
*Notes for Data Management*
# Due to Clouds, it might be worth declaring exposures 480637-480650 as bad.